OpenStudy (anonymous):

The Maryland Monkeys won six less than twice as many soccer games as The Pennsylvania Pigs. If the teams won a total of 48 games,how many games did each win ?

OpenStudy (anonymous):

The Maryland Monkeys\[48\cdot\dfrac{1}{2}-6=\] The Pennsylvania Pigs \[2\cdot18-6=\] See if you can see the logic.